A ball of mass 0.1kg is thrown vertically upwards with an initial velocity of 80m/s. Calculate the potential energy halfway up
1
Expert's answer
2021-08-16T17:43:36-0400

Let's calculate the maximum height

"H=\frac {v^2} {2g} =320m"

PE=mgh=0.1×10×320=320 J

at halfway up PE will be half of max = "\frac {320} {2} =160 J"

KE will be equal to PEmax=320 J